A place kicker kicks a football with a speed of 20.0 m/s at an angle of 40.0o upward from the horizontal. What is the speed of the ball after it has traveled a horizontal distance of 30.0 m?
Answer #1

U = 2om/sec Solution u 40 O 40. 3om U = U Cos (40) 20 cos (40) My= usin (40) Un = 15.3209 mlec 12.8558 m/sec At ang instant V2 لا : Hence at 30m. horizontal distance from, s= ut & se ut that Sx = Uut + 1 (0) tº 15-3209 (t) t = 1.95 81 sec. using k=ut